OpenStudy (anonymous):

Simplify the expression 144 to the negative one half power.

OpenStudy (ash2326):

We have \[\huge{144}^{\frac{-1}{2}}\] We know that \[a^{-b}=\frac{1}{a^b}\] Do you get this?

OpenStudy (anonymous):

sure....

OpenStudy (ash2326):

so we have \[\Huge \frac{1}{144^{\frac {1}{2}}}\] also \[a^\frac {1}{2}=\sqrt a\] can you simplify this now?

OpenStudy (anonymous):

1/12?

OpenStudy (ash2326):

yes, it's right

OpenStudy (ash2326):

do you understand?

OpenStudy (anonymous):

yep